OpenStudy (blackbird02):

What is the exact value of cos (7pi/6 + pi/4) ?

OpenStudy (blackbird02):

\[\cos (\frac{ 7\pi }{ 6 }) +\frac{ \pi }{ 4 }\]

OpenStudy (butterflydreamer):

Shouldn't you be using your cos sum rules? Like... cos (A +B) = cosAcosB - sinAsinB ??

OpenStudy (blackbird02):

i just want to know what is the equivalent angle of (7π/6) ? i've already solved this but i'm not that sure.

OpenStudy (mathmate):



OpenStudy (mathmate):